Estimated demand function for processed pork is Q=171-20p+20p_b+3p_c+2Y
Using the estimated demand function for processed pork in Namibia, show how the quantity
demanded at a given price changes as per capita income, Y, increases by N$100 a year.
[6 marks]
If the supply of corn by the Namibia is Q_n=a+bp and the supply by the rest of the world is Q_r=c+ep, what is the world supply? What is the world inverse supply? [6 marks]
The demand function for rose is Q=a-bp, and the supply function is Q=c+ep+ft, where a, b, c, e, and f are positive constants and t is the average temperature in a month. Show how the equilibrium quantity and price vary with temperature? [8 marks]

Question Two [10 marks]
What is the effect of a N$1 specific tax on equilibrium price and quantity if demand is perfectly inelastic? [4 marks]
The coconut oil demand function is Q=1200-9.5p+16.2p_p+0.2Y. Assume that p is initially N$0.45 per kg, p_p=N$0.31 per kg and Q=1275 thousand metri
1
Expert's answer
2015-06-01T10:34:38-0400
1. Q=171-20p+20pb+3pc+2y
If per capital income, Y, increases by N100, the quantity demanded at a given price will increase by 2*100 = 200 units.
2. If the supply of corn is Qn= a + bp and the supply by the rest of the word is¬ Qr=c+ep, the world supply is Qw = Qn + Qr = a + c + (b + e)*p.
3. Qd = a¬p-b, Qs = c+ep+ft. In equilibrium Qd = Qs, so: 
a¬pb = c+ep+ft
p*(e + b) = a - c - ft
p = (a - c - ft)/(e + b)
So, equilibrium price increases, when the temperature falls and equilibrium quantity falls too.
Question Two
If demand is perfectly inelastic, there will be no effect of a N$1 specific tax on equilibrium quantity, but equilibrium price will rise by N$1.